A tire of radius 1.5 ft rolls without slipping along a horizontal track with uniform angular acceleration of 3 rad/s^2 from rest.


(a) How long will it take the point on tire touching the track to acquire an angular frequency of 30 rad/s and how far horizontally could the tire travel at this instant?


(b) What was the tangential speed, centripetal acceleration, and tangential acceleration of the point on tire in part (a)?



1
Expert's answer
2021-12-03T12:44:40-0500

(a) The time is


"t=\frac \omega\alpha=10\text{ s}."


(b) At 30 rad/s, the tangential speed, centripetal acceleration, and tangential acceleration, are respectively:


"v=\omega r=45\text{ ft/s}.\\\space\\
a_c=\omega^2r=1350\text{ ft/s}^2,\\\space\\
a_\tau=\alpha r=4.5\text{ ft/s}^2."